A 20kg box is pushed across a flat floor with a force of 50N, if the

coefficient of kinetic friction between the floor and box is u = 0.20, what is

the acceleration of the box across the floor?

2.5 m/s/s

O 1.25 m/s/s

O 0.5 m/s/s

O m/s/s


1
Expert's answer
2021-01-30T06:26:23-0500
"ma=F-\mu mg\\20a=50-(0.2)(20)(9.8)\\a=0.54\frac{m}{s^2}"